# Intelligence Foundation Model: A New Perspective to Approach Artificial General Intelligence

### v2

@@ II-B Neuron Output Prediction
Every manifestation of intelligent behavior, whether it involves pattern recognition, memory formation, causal reasoning, or decision-making, emerges from the transformation of neuronal inputs into outputs through complex and adaptive neural dynamics. Inspired by this principle, we define a neuron output prediction objective for IFM, enabling it to learn from intelligent behaviors represented as neuronal input–output sequences. This objective unifies diverse forms of intelligent behavior within a single learning framework, allowing IFM to capture the general mechanisms underlying intelligence.
Formal Definition: With ground truth input stimulus sequence $(\mathbf{x}^{1},\mathbf{x}^{2},…,\mathbf{x}^{T})$ and the response sequence $(\mathbf{y}^{1},\mathbf{y}^{2},…,\mathbf{y}^{T})$ sampled from real-world human intelligent behaviors, the formulation of the neuron output prediction is defined as $\mathcal{L}=\min_{\mathcal{G}}\sum_{t=1}^{T}\ell(\mathbf{y}^{t},\hat{\mathbf{y}}^{t})$, with $\hat{\mathbf{y}}^{t}=\mathcal{G}(\mathbf{x}^{t},(\mathbf{s}^{t-1},\mathbf{e}^{t-1}))$. In the objective, $\mathcal{L}$ represents the total loss accumulated over a sequence of $T$ time steps, and $\ell(\mathbf{y}^{t},\hat{\mathbf{y}}^{t})$ measures the discrepancy between the real neuronal output $\mathbf{y}^{t}$ and predicted output by IFM $\hat{\mathbf{y}}^{t}$ at time $t$. By minimizing $\mathcal{L}$, IFM learns to predict correct neuronal outputs over time, capturing the temporal patterns and dependencies underlying intelligent behavior.
Concept Illustration: We illustrate the neuron output prediction concept using Pavlov’s classical conditioning experiment [pavlov], which is a foundational example of biological learning behavior [learning]. We first sample neuronal input–output sequences from the entire conditioning process, and then briefly explain how predicting neuronal outputs enables a state neural network to capture this biological-like learning behavior.
As illustrated in Fig. 4, the experiment shows that a dog learns to associate the sound of a ringing bell with the presentation of food through repeated training, eventually salivating at the sound alone. For simplicity, we model the dog with three functional neurons: a vision neuron that detects food ($F$), a sound neuron that perceives ringing ($R$), and a salivation neuron that triggers salivation ($S$). The overall conditioning process has three stages:
@@
Therefore, the entire conditioning process above can be represented as the temporally ordered sequences of neuronal inputs and outputs across the three stages combined, i.e., $\mathbf{x}=\{(F,\neg R),(\neg F,R),(F,R),(F,R),(\neg F,R)\}$ and $\mathbf{y}=\{S,\neg S,S,S,S\}$. More neuronal sequences can be obtained by varying the trial patterns, such as increasing or decreasing the number of trials in each stage, to enrich the dataset and improve sample diversity.
Under the objective of neuron output prediction, a state neural network is expected to be trained to generate proper $\mathbf{y}$ given $\mathbf{x}$, indicating the possession of the conditioning-like learning ability. Specifically, we train the state neural network $\mathcal{G}$ using the above neuronal input–output sequences, where two input neurons receive $\mathbf{x}$ and one output neuron generates $\mathbf{y}$. The network includes multiple hidden neurons to learn the mapping from $\mathbf{x}$ to $\mathbf{y}$ by minimizing the neuron output prediction loss $\mathcal{L}$. Through this training process, $\mathcal{G}$ adjusts its internal parameters that regulate neuron function, neuron connectivity, and neuron plasticity to fit the conditioning mechanism and reproduces similar neuronal output patterns during inference. Note that the trained network is not intended to memorize specific training samples, but rather to learn the fundamental principles of classical conditioning that underlie a large number of training samples.
Figure 4: Pavlov’s classical conditioning experiment.
@@ III IFM Training
To train IFM under the neuron output prediction objective, a major challenge lies in constructing neuronal input–output datasets that capture the temporal evolution of neural activity underlying intelligent behaviors. Since IFM is designed to learn the time-dependent mapping between neuronal inputs and outputs, its training data should ideally represent the dynamic transformation of signals as the neural system interacts with its environment. Each data sample consists of a pair of neuronal input–output sequences, expressed as $d=\{(\mathbf{x}^{1},...,\mathbf{x}^{T}),(\mathbf{y}^{1},...,\mathbf{y}^{T})\}$, while the overall dataset comprises multiple such samples, denoted by $D=\{d_{1},d_{2},\ldots,d_{n}\}$. Following the paradigm of FMs, data samples should be collected from a wide range of individuals engaged in diverse intelligent behaviors. Such data samples can be obtained in two ways:
@@
  • Direct Neuronal Sampling: This approach acquires neuronal input–output data directly from biological brains by recording synaptic activity across neurons as subjects perform intelligent behaviors. It provides the most biologically faithful supervision for IFM, as it reflects the genuine neuronal computations underlying cognition. However, large-scale data collection remains technically infeasible because capturing the full dynamics of even small cortical circuits at cellular resolution remains an open challenge. Current techniques, such as Neuropixels probes [Neuropixels] and calcium imaging [ultrasensitive], can only record limited subsets of neurons and modalities, making this method suitable primarily for small-scale or proof-of-concept IFMs.
@@
  • Indirect Neuronal Sampling: This scalable alternative captures functional equivalents of neuronal inputs and outputs through fully instrumented, human-in-the-loop embodiments, such as a neurosensory interface (a wearable multimodal sensing system) or an avatar (a digital or robotic agent with human-like sensing and actuation). These embodiments perceive environmental stimuli (visual, auditory, tactile, proprioceptive) and record observable behavioral responses (speech, gestures, motor actions). The sensory streams serve as proxies for neuronal inputs, while behavioral responses represent neuronal outputs. Recording paired temporal sequences during natural interactions produces high-dimensional data encoding the dynamics of intelligent behavior. Training IFM on such data enables it to learn how sensory experiences evolve into actions over time, effectively capturing neuron-like transformation principles without requiring direct observation of biological neurons.
@@
As neural recording technologies progress, direct and indirect neuronal sampling are expected to converge, enabling IFM to move beyond emulating intelligent behavior toward internalizing its underlying mechanisms, thereby closing the loop between cognition and computation. With these data samples, IFM implemented by differentiable state neural networks can be efficiently trained via backpropagation. Given the streaming nature of the training data, variants of Truncated Backpropagation Through Time (TBPTT) [tbptt] can be employed to reduce computational complexity. We present a toy example in which IFM is used to train an artificial Pong player, demonstrating the overall design process available at https://github.com/brcai/IFM_pong.
IV Discussion and Outlook
@@
Discussion: Although humans differ across age, culture, and individual biology, intelligence consistently arises from shared fundamental principles embedded in the brain. Building on this premise, IFM leverages the foundation model paradigm to capture these underlying mechanisms directly from neuronal dynamics. This approach defines a fundamentally new pathway toward AGI, one that models the mechanisms of intelligence itself rather than its fragmented expressions. We envisage a pragmatic and scalable roadmap toward AGI through IFM from two complementary perspectives:
@@
  • The first trajectory focuses on biological scaling, starting from simpler neural systems or lower organisms and progressively advancing to more complex brains, to systematically learn the principles of intelligence across increasing levels of complexity;
@@
  • The second trajectory focuses on functional scaling, developing IFMs for specific robotic applications such as industrial manipulators, household assistants or autonomous agents and gradually integrating these specialized capabilities into a unified, general-purpose intelligent system.
By integrating these two trajectories, IFM offers a stepwise, interpretable, and biologically inspired pathway toward AGI.
Outlook: A key implication of IFM is its potential in conceptual unification of human intelligence and machine intelligence. By modeling the universal principles that generate intelligence, IFM’s artificial substrates (state neural networks) can replicate the functional dynamics of human cognition without demanding any specific biological neuron cells. The success of IFM will show that consciousness, learning, reasoning, and memory emerge from dynamic, recurrent, and plastic interactions, whether instantiated in biological or artificial media. This opens the door to transformative possibilities, such as continuous mind uploading [slow], where biological neurons are gradually replaced or augmented with artificial state neurons while preserving the continuity of consciousness. Such an approach safeguards personal identity and cognitive continuity, surpassing traditional “copy-based” mind uploading paradigms [upload], and reframes our understanding of intelligence as substrate-independent. That paints a transformative vision for the future of intelligent life, where the boundary between natural and artificial intelligence becomes increasingly seamless and where human-level cognition can be instantiated beyond the biological brain.
IFM envisions a future where intelligence is no longer confined to biology. Humans and machines could share a common cognitive substrate, enabling seamless augmentation, collaboration, and extension of minds. Adaptive artificial neurons could integrate with biological networks, expanding memory, perception, and reasoning capacities beyond natural limits. In this new paradigm, consciousness, learning, and creativity are universal phenomena that can be instantiated in multiple substrates, offering humanity unprecedented opportunities to explore, understand, and even evolve intelligence itself. IFM does not merely aim to replicate intelligence, but it lays the foundation for a future in which the boundaries between human, machine, and mind are dynamically intertwined, ushering in a new era of cognitive life.

### v3

@@ II-B Neuron Output Prediction
Every manifestation of intelligent behavior, whether it involves pattern recognition, memory formation, causal reasoning, or decision-making, emerges from the transformation of neuronal inputs into outputs through complex and adaptive neural dynamics. Based on this principle, we introduce a neuron output prediction objective for IFM to learn from intelligent behavior represented as neuronal input–output sequences. This objective unifies diverse intelligent behaviors within a single learning framework and enables IFM to capture the underlying principles of intelligence.
Formal Definition: With ground truth input stimulus sequence $(\mathbf{x}^{1},\mathbf{x}^{2},…,\mathbf{x}^{T})$ and the response sequence $(\mathbf{y}^{1},\mathbf{y}^{2},…,\mathbf{y}^{T})$ sampled from real-world human intelligent behavior, the formulation of the neuron output prediction is defined as $\mathcal{L}=\min_{\mathcal{G}}\sum_{t=1}^{T}\ell(\mathbf{y}^{t},\hat{\mathbf{y}}^{t})$, with $\hat{\mathbf{y}}^{t}=\mathcal{G}(\mathbf{x}^{t},(\mathbf{s}^{t-1},\mathbf{e}^{t-1}))$. In the objective, $\mathcal{L}$ represents the total loss accumulated over a sequence of $T$ time steps, and $\ell(\mathbf{y}^{t},\hat{\mathbf{y}}^{t})$ measures the discrepancy between the real neuronal output $\mathbf{y}^{t}$ and predicted output by IFM $\hat{\mathbf{y}}^{t}$ at time $t$. By minimizing $\mathcal{L}$, IFM learns to predict correct neuronal outputs over time, thereby capturing the temporal patterns and dependencies underlying intelligent behavior.
Figure 4: Pavlov’s classical conditioning experiment.
Concept Illustration: We illustrate the neuron output prediction concept using Pavlov’s classical conditioning experiment [pavlov], which is a canonical example of biological learning behavior [learning]. We first sample neuronal input–output sequences from the entire conditioning process, then show how predicting these neuronal outputs allows a state neural network to emulate this biological learning behavior.
As illustrated in Fig. 4, the experiment shows that, through repeated training, a dog learns to associate the sound of a ringing bell with the presentation of food, eventually salivating at the sound alone. For simplicity, we represent the dog with three functional neurons: a vision neuron that detects food ($F$), a sound neuron that perceives ringing ($R$), and a salivation neuron that triggers salivation ($S$). The overall conditioning process has three stages:
@@
Therefore, the entire conditioning behavior can be represented as temporal sequences of neuronal inputs and outputs spanning the three stages, i.e., $\mathbf{x}=\{(F,\neg R),(\neg F,R),(F,R),(F,R),(\neg F,R)\}$ and $\mathbf{y}=\{S,\neg S,S,S,S\}$. Additional sequences can be generated by varying trial patterns, e.g., changing the number of trials per stage, to enrich the dataset and increase sample diversity.
Under the objective of neuron output prediction, a state neural network is trained to predict $\mathbf{y}$ from $\mathbf{x}$, demonstrating conditioning-like learning ability. Specifically, we train the state neural network $\mathcal{G}$ using the above neuronal input–output sequences, where two input neurons receive $\mathbf{x}$ and one output neuron generates $\mathbf{y}$. The network includes multiple hidden neurons to learn the mapping from $\mathbf{x}$ to $\mathbf{y}$ by minimizing the neuron output prediction loss $\mathcal{L}$. Through this training process, $\mathcal{G}$ adjusts its internal parameters that regulate neuron function, neuron connectivity, and neuronal plasticity to emulate the conditioning mechanism and reproduce corresponding neuronal output patterns. Importantly, the network is not meant to memorize individual trials, but to internalize the fundamental principles of classical conditioning that generalize across many samples.
@@ III IFM Training
A major challenge in training IFM under the neuron output prediction objective is constructing neuronal input–output datasets that capture the temporal evolution of neural activities underlying intelligent behavior. Because IFM learns the time-dependent mapping between neuronal inputs and outputs, the training data should reflect the dynamic transformation of signals as the neural system interacts with its environment. Each data sample consists of a pair of neuronal input–output sequences, expressed as $d=\{(\mathbf{x}^{1},...,\mathbf{x}^{T}),(\mathbf{y}^{1},...,\mathbf{y}^{T})\}$, while the overall dataset comprises multiple such samples, denoted by $D=\{d_{1},d_{2},\ldots,d_{n}\}$. In line with the FM paradigm, data samples should be collected from a broad range of individuals performing diverse intelligent behaviors. Such data samples can be obtained in two ways:
@@
  • Direct Neuronal Sampling: This approach acquires neuronal input–output data directly from biological brains by recording synaptic activity across neurons as subjects perform intelligent behavior. It provides the most biologically faithful supervision for IFM, as it reflects the genuine neuronal computations underlying cognition. However, large-scale data collection is technically infeasible, as capturing the full dynamics of even small cortical circuits at cellular resolution remains an open challenge. Current techniques, such as Neuropixels probes [Neuropixels] and calcium imaging [ultrasensitive], can only record limited subsets of neurons and modalities, making this method suitable primarily for small-scale or proof-of-concept IFMs.
@@
  • Indirect Neuronal Sampling: This scalable alternative captures functional equivalents of neuronal inputs and outputs via fully instrumented, human-in-the-loop embodiments, such as neurosensory interfaces (wearable multimodal sensing systems) or avatars (digital or robotic agents with human-like perception and actuation). These embodiments perceive environmental stimuli (visual, auditory, tactile, proprioceptive) and record observable behavioral responses (speech, gestures, motor actions). The sensory streams serve as proxies for neuronal inputs, while behavioral responses represent neuronal outputs. Recording paired temporal sequences during natural interactions produces high-dimensional data encoding the dynamics of intelligent behavior. Training IFM on such data enables it to learn how sensory experiences evolve into actions over time, effectively capturing neuron-like transformation principles without requiring direct observation of biological neurons.
@@
With advances in neural recording technologies, direct and indirect neuronal sampling are expected to converge. This convergence will enable IFM to go beyond emulating intelligent behavior, moving toward internalizing its underlying principles and thus closing the loop between cognition and computation. Using these data samples, IFM implemented by differentiable state neural networks can be efficiently trained via backpropagation. Given the streaming nature of the training data, variants of Truncated Backpropagation Through Time (TBPTT) [tbptt] can be employed to reduce computational complexity. We present a toy example in which IFM is used to train an artificial Pong player, demonstrating the overall design process available at https://github.com/brcai/IFM_pong.
IV Discussion and Significance
@@
Discussion: Despite variations in age, culture, and individual biology, human intelligence consistently arises from shared fundamental principles embedded in the brain. Building on this premise, IFM leverages the foundation model paradigm to capture these underlying mechanisms of neuronal dynamics. This approach defines a fundamentally new pathway toward AGI, one that models the principles of intelligence itself rather than its fragmented expressions. We envisage a pragmatic and scalable roadmap toward AGI through IFM from two complementary trajectories:
@@
  • Biological scaling begins with simpler neural systems or lower organisms and progressively advances to more complex brains, systematically revealing the principles of intelligence across increasing levels of complexity.
@@
  • Functional scaling develops IFMs for specific robotic applications, such as industrial manipulators or household assistants, and gradually integrates these specialized capabilities into a unified, general-purpose intelligent system.
Despite its promise, realizing this vision faces several major challenges. Acquiring high-resolution, longitudinal neural activity data across diverse species and brain regions is technically demanding. Designing state neural networks that faithfully capture recurrent, plastic, and temporally extended neural dynamics while remaining computationally tractable presents significant modeling challenges. Finally, training IFM to reproduce complex neural dynamics across multiple systems and scales requires novel optimization strategies to ensure convergence and generalization. Overcoming these challenges will be essential to advancing IFM as a practically viable approach toward general-purpose intelligence.
Significance: By reframing intelligence as the product of underlying structural neural dynamics rather than isolated behavioral manifestations, IFM provides a paradigm shift with broad implications for both neuroscience and AI. This perspective suggests that AGI may not require reproducing the full biological intricacies of the brain, but instead capturing the dynamical principles that allow artificial neural systems to generate diverse intelligent behavior. IFM thus offers a unifying framework that bridges three traditionally separate areas: neural circuit dynamics, models of learning and adaptive behavior, and large-scale AI systems. By grounding artificial systems in these shared principles, IFM opens a pathway toward constructing models whose generalization capabilities arise from neuronal structure rather than model scale alone. At the same time, IFM highlights critical unresolved questions: What minimal neuron functions are sufficient to support general-purpose intelligence? Which aspects of biological plasticity are essential for adaptive behavior? How can artificial systems reconcile the tension between biological complexity and computational tractability? Addressing these questions with IFM could not only advance AGI but also deepen our understanding of intelligence as a natural phenomenon.
Ultimately, IFM provides more than a technical proposal; it outlines a long-term scientific agenda. By integrating biological and functional scaling, IFM charts a stepwise, interpretable, and experimentally grounded path toward general-purpose intelligent systems. Its challenges are substantial, but so too is its potential to reshape how we study, model, and engineer intelligence across biological and artificial domains.